OpenStudy (mathhelp346):

Suppose f(x)=ax^4+bx^2+x+5. Find f(4) given f(-4)=3

geerky42 (geerky42):

Let's just plug in \(-4:~f(-4)=4^4a+4^2b-4+5 = 3\) Let \(u=4^4a+4^2b,\) we have \(f(-4)=u+1=3\) So \(\underline{u = 2}\) Now plug in \(4\), we have \[\begin{array}~f(4) &= 4^4a+4^2b+4+5\\~\\ &= u+4+5\\~\\&=2+4+5\\~\\&=\boxed{11}\end{array}\]

OpenStudy (mathhelp346):

What's u?

geerky42 (geerky42):

Just placehold for \(4^4a+4^2b\).

OpenStudy (mathhelp346):

if you plug in -4, wouldn't it be \[-4a^4-4b^2 \] instead of \[4^4a+4^2b\]

geerky42 (geerky42):

Actually, it would be \((-4)^4a+(-4)^2b\) which is equal to \(4^4a+4^2b\) because even exponent cancels minus sign.

geerky42 (geerky42):

Do you understand what I did? Ask question if you are confused/

OpenStudy (mathhelp346):

Yes, thank you so much!